# Improved Regret of Linear Ensemble Sampling

### v1

@@ Abstract
In this work, we close the fundamental gap of theory and practice by providing an improved regret bound for linear ensemble sampling. We prove that with an ensemble size logarithmic in $T$, linear ensemble sampling can achieve a frequentist regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$, matching state-of-the-art results for randomized linear bandit algorithms, where $d$ and $T$ are the dimension of the parameter and the time horizon respectively. Our approach introduces a general regret analysis framework for linear bandit algorithms. Additionally, we reveal a significant relationship between linear ensemble sampling and Linear Perturbed-History Exploration (LinPHE), showing that LinPHE is a special case of linear ensemble sampling when the ensemble size equals $T$. This insight allows us to derive a new regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE, independent of the number of arms. Our contributions advance the theoretical foundation of ensemble sampling, bringing its regret bounds in line with the best known bounds for other randomized exploration algorithms.
@@ 1 Introduction
In this work, we close this fundamental gap by providing an improved regret bound for linear ensemble sampling. We prove that linear ensemble sampling with an ensemble size logarithmic in $T$ can still attain a frequentist regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$, marking the first time that linear ensemble sampling achieves a state-of-the-art result for randomized linear bandit algorithms. Our approach not only improves upon the regret bound but also simplifies the algorithm by avoiding the use of symmetrized perturbations, making it more practical for implementation. For regret analysis, we present a general, concise framework for analyzing linear bandit algorithms, which may be of independent interest. Furthermore, we rigorously reveal the significant relationship between ensemble sampling and PHE for the first time, showing that in the regime where the ensemble size equals $T$, linear PHE (LinPHE) is a special case of linear ensemble sampling. With this new insight, we can use the regret analysis for ensemble sampling to derive a new regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE, which is independent of the number of arms $K$.
@@ 3 Ensemble Sampling for Linear Bandits
Ensemble sampling is capable of being generalized to complex settings whenever solving minimization problem (1) is tractable. Especially when incremental updates of the minimization problem are cheap, for instance with neural networks or other gradient descent-based models, ensemble sampling can be an efficient exploration strategy. The algorithm may simply store an ensemble of models, sample one to select an action, and then update the models incrementally based on the observed reward and generated perturbation.
@@ 4 Regret Bound of Linear Ensemble Sampling
As in Lu and Van Roy [16] and Qin et al. [21], we study the finite-armed problem setting. Both studies analyze the excess regret of ensemble sampling compared to Thompson sampling through an information theoretical approach. However, direct comparisons of the regret bounds are non-trivial. The analysis by Lu and Van Roy [16] includes an error admitted by the authors and Qin et al. [21] analyze the Bayesian regret, which is a weaker notion of regret than the frequentist regret that we analyze in this work. Along with Janz et al. [10], our result also makes a progress in reducing the size of the ensemble compared to  Lu and Van Roy [16] and Qin et al. [21]. The size of the ensemble required by Janz et al. [10] is $\Theta(d\log T)$.This requirement implies that their ensemble size may be smaller than ours when $K$ is larger than $d$ and also allows $K$ to be infinite. However, it is important to note that their resulting regret bound of $\widetilde{\mathcal{O}}(d^{5/2}\sqrt{T})$ is clearly sub-optimal compared to regret bounds of other randomized exploration algorithms. Theorem 1 achieves the tighter regret bound while simultaneously reducing the size of the ensemble.
@@ 5.1 General Regret Analysis for Linear Bandits
The audience well-versed in the regret analysis of randomized algorithms such as TS and PHE would recognize that bounding the regret using the probability of being optimistic is a standard procedure, also presented in Theorem 1 of Kveton et al. [12] and Theorem 2 of Janz et al. [10] as generalizations of the results in Agrawal and Goyal [4] and Abeille and Lazaric [2] respectively. However, our regret analysis offers much more concise approach than the existing techniques, which can be of independent interest beyond the analysis of ensemble sampling.
Theorem 2 states that if the two conditions, specifically concentration in (2) and optimism in (3), are met, then the algorithm achieves $\sqrt{T}$ regret. We provide the proof of Theorem 2 in Appendix B. Our proof technique generalizes the well-studied analysis of Abeille and Lazaric [2]. While their work poses conditions on a $d$-dimensional perturbation vector that is added to the ridge estimator, we do not assume the use of ridge regression nor we assume that the estimator is perturbed. Instead, we only pose conditions on the final estimator the algorithm exploits. Due to this generalization, Theorem 2 is even capable of inducing the regret bound of LinUCB [1], which always opts for an optimistic estimator, by setting $\gamma=2\beta_{T}$ and $p=1$ with appropriate concentration events assigned to $\mathcal{E}_{1,t}$ and $\mathcal{E}_{2,t}$. In addition, there are several improvements that simplify the proof which are worth noting. To exploit the optimism condition, we apply Markov’s inequality on a well-defined random variable. The proof of Abeille and Lazaric [2] relies on defining a conditional distribution, conditioned on both history and the event of being optimistic. However, such distribution may not be well-defined if the probability of the event is 00 for given history. Janz et al. [10] try to solve this problem by separately handling such exceptional cases using conditional measures. However, their conditional measures depend on the random history, leading the probability $p$ to be a random variable, which complicates the analysis. We also note that our proof does not require convex analysis studied in Abeille and Lazaric [2].
@@ 5.2 Proof of Regret Bound in Theorem 1
We prove the regret bound of linear ensemble sampling stated in Theorem 1. To apply Theorem 2, high-probabilities of the sequences of events, namely $\{\mathcal{E}_{1,t},\mathcal{E}_{2,t}\}_{t=1}^{T}$, should be guaranteed with an appropriate values of $\gamma$ and $p$. We show that separate constraints can be imposed on the randomness of the rewards and the perturbations respectively to guarantee the probabilities of the events. We begin by decomposing the estimator into two parts: one that fits the observed rewards and the other that perturbs the estimator.
@@
We present a straightforward proof of Lemma 2 in Appendix C. We significantly deviate from the analyses of Abeille and Lazaric [2] and Janz et al. [10] in the method of guaranteeing the optimism condition. Their analyses require a $d$-dimensional perturbation vector to have a constant probability of having positive component, so-called anti-concentrated, in “every” possible direction in $\mathbb{R}^{d}$ since they only prove the existence of a direction that implies optimism. We observe and exploit the fact that it suffices to consider just “one” direction, specifically $U_{t-1}$, to produce an optimistic estimator. Since $U_{t-1}$ depends only on the sequence of selected arms, the dependency between $U_{t-1}$ and $\mathbf{Z}_{t}$ decouples when the dependency between $\{X_{t}\}_{t=1}^{T}$ and $\{\mathbf{Z}_{t}\}_{t=1}^{T}$ are decoupled, which we later achieve by taking the union bound in a unique way.
@@
where Lemma 2 bridges the anti-concentration on the left hand side to the optimism on the right hand side. This implication relationship is converted to the following probability inequality:
@@
By the definition of ~$\widetilde{\mathcal{E}}_{2,t}$, the right hand side is bounded below by $p$, implying optimism condition (3). The probability of failure is bounded by the union bound and Lemma 1. ∎
@@
Suppose Algorithm 1 is run with parameters specified in Theorem 1. Fix $t\in[T]$ and $j\in[m]$. Suppose the sequence of arms $X_{1},\ldots,X_{t}$ is chosen arbitrarily randomly, not necessarily by the the agent. Let $\widetilde{\theta}_{t}^{j}$ be defined as in Eq. (5). Then, with probability at least $1-\delta/T$, $\|\widetilde{\theta}_{t}^{j}\|_{V_{t}}\leq\widetilde{\gamma}_{T}$ holds.
The following fact describes an anti-concentration property of Gaussian distribution, which follows from the fact that a linear combination of independent Gaussians is again Gaussian.
@@
All the building blocks we need to prove Theorem 1 is ready. The proof illustrates that the events specified in Lemma 3 occur with high probability.
@@
Since we have verified that the expectation of the right hand side is greater than $p_{N}/2$, Azuma-Hoeffding inequality (Lemma 12) implies that $\mathbb{P}(\frac{1}{m}\sum_{j=1}^{m}I^{j}_{t}<p_{N}/4)\leq\exp\left(-p_{N}^{2}m/8\right)$. Recall that this result is obtained assuming that $X_{1},\ldots,X_{T}$ are fixed. We take the union bound over all possible sequences of arms. However, a naïve union bound multiplies $K^{T}$ to the failure probability, which leads to an undesirable result of $m$ scaling linearly with $T$. We present the following claim inspired by an observation from Lu and Van Roy [16] that a permutation of selected arms can be regarded as equivalent. We note that the strong result of Lemma 6 in Lu and Van Roy [16] is not applicable to our setting since we do not assume that $\{\eta_{t}\}_{t=1}^{T}$ is distributed identically nor independently. Although we present all the main ideas to support Claim 1 in this section, there may be a few points that readers find require further justification. Due to limited space, we provide a full, rigorous justification of Claim 1 in Appendix F, where we present a different perspective on the sampling of perturbation. We note that this claim is not a conjecture nor an assumption.
Claim 1.
@@
The key observation in Claim 1 is that the perturbation vector consists of i.i.d. components, hence its distribution is invariant under independent permutations. Therefore, the distributions of $\widetilde{\theta}_{t-1}^{j}$ and $U_{t-1}^{\top}{\mathbf{Z}}_{t}^{j}$ remain invariant under the permutation of selected arms. Although the sequence of arms and the perturbation vector are not independent as a whole, the permutation that sorts the selected arms preserves the distribution of $\mathbf{Z}_{t}$ since $X_{t}$ and $Z_{t}$ are independent for all $t\in[T]$. The number of equivalence classes up to permutation over sequences of arms with lengths at most $T-1$ is less than $T^{K}$, since each arm can be selected 00 to $T-1$ times inclusively. Therefore, we take the union bound over the $T^{K}$ sequences of arms and attain $\mathcal{E}_{2}^{*}$.
@@ 6 Ensemble Sampling and Perturbed-History Exploration
LinPHE is linear ensemble sampling with $T$ models: Conversely, since LinPHE can be regarded as linear ensemble sampling with $T$ models, it is possible to derive a regret bound of LinPHE by following the proof of Theorem 1. We present Corollary 1, which states a new regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE. Note that in this case, the regret bound is independent of $K$.
@@
Kveton et al. [12] provide a $\widetilde{\mathcal{O}}(d\sqrt{T\log K})$ regret bound when the number of arms is finite. Our result is the first to prove that LinPHE achieves a $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret bound that is independent of the number of arms. Hence, we view our overall results as a generalization of Kveton et al. [12]. It is widely observed that assuming the size of the arm set to be $K$ may lead to an interchanging of a $\sqrt{d}$ factor with a $\sqrt{\log K}$ factor in the regret bound [4], although attaining such reduction may not always be done in a trivial manner [5, 8, 6]. Our focus is not merely on proving another regret bound for LinPHE, but rather on highlighting the close relationship between linear ensemble sampling and LinPHE, which, to our knowledge, has been overlooked in the literature.
The proof of Corollary 1 follows the proof of Theorem 1. Note that the latter part of the proof of Theorem 1 focuses on decoupling the dependency between $\{X_{t}\}_{t=1}^{T}$ and $\mathbf{Z}_{t}$. However, in the case of LinPHE, they are already independent since the perturbation sequence is freshly sampled at every time step, enabling a more elegant and concise proof. Especially, as it skips the parts that require the number of arms to be finite, for instance the use of Claim 1, Corollary 1 holds even when the number of arms is infinite. The whole proof is presented in Appendix G.
@@ Appendix A Notations
We summarizes the notations in this paper in Table 2 and Table 3
@@
$\mathcal{E}$ High probability event
@@ Appendix B Proof of Theorem 2
As alluded in the discussion of Theorem 2, we utilize Markov’s inequality and the random variable of interest is $X_{t}^{\top}\theta_{t}\mathds{1}\left\{\mathcal{E}_{1,t}\cap\mathcal{E}_{2,t}\right\}$. However, this random variable may not be non-negative, precluding the use of Markov’s inequality. The following lemma shows that adding an appropriate term guarantees its non-negativity.
@@
We lower-bound the probability on the left hand side utilizing condition (3). Suppose the event of interest in condition (3), namely $((x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t})\lor\mathcal{E}_{2,t}^{\mathsf{C}}$, holds. Under the event, either $\mathcal{E}_{2,t}^{\mathsf{C}}$ or $(x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t}\land\mathcal{E}_{2,t}$ holds. Under $\mathcal{E}_{2,t}^{\mathsf{C}}$, $g_{t}(\theta^{*},\mathcal{E}_{2,t})\leq g_{t}\left(\theta_{t},\mathcal{E}_{1,t}\cap\mathcal{E}_{2,t}\right)$ becomes $0\leq 0$, which trivially holds. Otherwise, we have $(x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t}\land\mathcal{E}_{2,t}$. Under this event, we have that
@@
The second sum is bounded by Azuma-Hoeffding inequality. Note that $0\leq\left\|X_{t}\right\|_{V_{t-1}^{-1}}\leq\sqrt{\lambda_{\max}(V_{t-1}^{-1})}\left\|X_{t}\right\|_{2}\leq\frac{1}{\sqrt{\lambda}}$, where $\lambda_{\max}(\cdot)$ denotes the maximum eigenvalue. By Lemma 12, with probability at least $1-\delta$, it holds that

### v3

@@ Abstract
In this work, we close the fundamental gap of theory and practice by providing an improved regret bound for linear ensemble sampling. We prove that with an ensemble size logarithmic in $T$, linear ensemble sampling can achieve a frequentist regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$, matching state-of-the-art results for randomized linear bandit algorithms, where $d$ and $T$ are the dimension of the parameter and the time horizon respectively. Our approach introduces a general regret analysis framework for linear bandit algorithms. Additionally, we reveal a significant relationship between linear ensemble sampling and Linear Perturbed-History Exploration (LinPHE), showing that LinPHE is a special case of linear ensemble sampling when the ensemble size equals $T$. This insight allows our analysis framework to derive a regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE, independent of the number of arms. Our techniques advance the theoretical foundation of ensemble sampling, bringing its regret bounds in line with the best known bounds for other randomized exploration algorithms.
@@ 1 Introduction
In this work, we close this fundamental gap by providing an improved regret bound for linear ensemble sampling. We prove that linear ensemble sampling with an ensemble size logarithmic in $T$ can still attain a frequentist regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$, marking the first time that linear ensemble sampling achieves a state-of-the-art result for randomized linear bandit algorithms. Our approach not only improves upon the regret bound but also simplifies the algorithm by avoiding the use of symmetrized perturbations, making it more practical for implementation. For regret analysis, we present a general, concise framework for analyzing linear bandit algorithms, which may be of independent interest. Furthermore, we rigorously reveal the significant relationship between ensemble sampling and PHE for the first time, showing that in the regime where the ensemble size equals $T$, linear PHE (LinPHE) is a special case of linear ensemble sampling. With this new insight, we can use the regret analysis for ensemble sampling to derive a regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE, which is independent of the number of arms $K$.
@@ 3 Ensemble Sampling for Linear Bandits
Ensemble sampling is capable of being generalized to complex settings whenever solving minimization problem (1) is tractable. Especially when incremental updates of the minimization problem are cheap, for instance, with neural networks or other gradient descent-based models, ensemble sampling can be an efficient exploration strategy. The algorithm may simply store an ensemble of models, sample one to select an action, and then update the models incrementally based on the observed reward and generated perturbation.
@@ 4 Regret Bound of Linear Ensemble Sampling
As in Lu and Van Roy [16] and Qin et al. [21], we study the finite-armed problem setting. Both studies analyze the excess regret of ensemble sampling compared to Thompson sampling through an information theoretical approach. However, direct comparisons of the regret bounds are non-trivial. The analysis by Lu and Van Roy [16] includes an error admitted by the authors and Qin et al. [21] analyze the Bayesian regret, which is a weaker notion of regret than the frequentist regret that we analyze in this work. Along with Janz et al. [10], our result also makes progress in reducing the size of the ensemble compared to  Lu and Van Roy [16] and Qin et al. [21]. The size of the ensemble required by Janz et al. [10] is $\Theta(d\log T)$. This requirement implies that their ensemble size may be smaller than ours when $K$ is larger than $d$ and also allows $K$ to be infinite. However, it is important to note that their resulting regret bound of $\widetilde{\mathcal{O}}(d^{5/2}\sqrt{T})$ is clearly sub-optimal compared to regret bounds of other randomized exploration algorithms. Theorem 1 achieves the tighter regret bound while simultaneously reducing the size of the ensemble.
@@ 5.1 General Regret Analysis for Linear Bandits
The audience well-versed in the regret analysis of randomized algorithms such as TS and PHE would recognize that bounding the regret using the probability of being optimistic is a standard procedure, also presented in Theorem 1 of Kveton et al. [12] and Theorem 2 of Janz et al. [10] as generalizations of the results in Agrawal and Goyal [4] and Abeille and Lazaric [2] respectively. However, our regret analysis offers a much more concise approach than the existing techniques, which can be of independent interest beyond the analysis of ensemble sampling.
Theorem 2 states that if the two conditions, specifically concentration in (2) and optimism in (3), are met, then the algorithm achieves $\sqrt{T}$ regret. We provide the proof of Theorem 2 in Appendix B. Our proof technique generalizes the well-studied analysis of Abeille and Lazaric [2]. While their work poses conditions on a $d$-dimensional perturbation vector that is added to the ridge estimator, we do not assume the use of ridge regression or assume that the estimator is perturbed. Instead, we only pose conditions on the final estimator the algorithm exploits. Due to this generalization, Theorem 2 is even capable of inducing the regret bound of LinUCB [1], which always opts for an optimistic estimator, by setting $\gamma=2\beta_{T}$ and $p=1$ with appropriate concentration events assigned to $\mathcal{E}_{1,t}$ and $\mathcal{E}_{2,t}$. In addition, there are several improvements that simplify the proof which are worth noting. To exploit the optimism condition, we apply Markov’s inequality to a well-defined random variable. This allows us to avoid using conditional expectations conditioned on the optimism event, which requires additional handling when the event has probability zero. Furthermore, the theorem and its proof explicitly specify the high-probability event under which the analysis is conducted. We also note that our proof does not require convex analysis studied in Abeille and Lazaric [2].
@@ 5.2 Proof of Regret Bound in Theorem 1
We prove the regret bound of linear ensemble sampling stated in Theorem 1. To apply Theorem 2, high probabilities of the sequences of events, namely $\{\mathcal{E}_{1,t},\mathcal{E}_{2,t}\}_{t=1}^{T}$, should be guaranteed with appropriate values of $\gamma$ and $p$. We show that separate constraints can be imposed on the randomness of the rewards and the perturbations respectively to guarantee the probabilities of the events. We begin by decomposing the estimator into two parts: one that fits the observed rewards and the other that perturbs the estimator.
@@
We present a straightforward proof of Lemma 2 in Appendix C. We significantly deviate from the analyses of Abeille and Lazaric [2] and Janz et al. [10] in the method of guaranteeing the optimism condition. Their analyses require a $d$-dimensional perturbation vector to have a constant probability of having a positive component, so-called anti-concentrated, in “every” possible direction in $\mathbb{R}^{d}$ since they only prove the existence of a direction that implies optimism. We observe and exploit the fact that it suffices to consider just “one” direction, specifically $U_{t-1}$, to produce an optimistic estimator. Since $U_{t-1}$ depends only on the sequence of selected arms, the dependency between $U_{t-1}$ and $\mathbf{Z}_{t}$ decouples when the dependency between $\{X_{t}\}_{t=1}^{T}$ and $\{\mathbf{Z}_{t}\}_{t=1}^{T}$ is decoupled, which we later achieve by taking the union bound in a unique way.
@@
where Lemma 2 bridges the anti-concentration on the left-hand side to the optimism on the right-hand side. This implication relationship is converted to the following probability inequality:
@@
By the definition of ~$\widetilde{\mathcal{E}}_{2,t}$, the right-hand side is bounded below by $p$, implying optimism condition (3). The probability of failure is bounded by the union bound and Lemma 1. ∎
@@
Suppose Algorithm 1 is run with parameters specified in Theorem 1. Fix $t\in[T]$ and $j\in[m]$. Suppose the sequence of arms $X_{1},\ldots,X_{t}$ is chosen arbitrarily randomly, not necessarily by the agent. Let $\widetilde{\theta}_{t}^{j}$ be defined as in Eq. (5). Then, with probability at least $1-\delta/T$, $\|\widetilde{\theta}_{t}^{j}\|_{V_{t}}\leq\widetilde{\gamma}_{T}$ holds.
The following fact describes an anti-concentration property of the Gaussian distribution, which follows from the fact that a linear combination of independent Gaussians is again Gaussian.
@@
All the building blocks we need to prove Theorem 1 are ready. The proof illustrates that the events specified in Lemma 3 occur with high probability.
@@
Since we have verified that the expectation of the right-hand side is greater than $p_{N}/2$, the Azuma-Hoeffding inequality (Lemma 12) implies that $\mathbb{P}(\frac{1}{m}\sum_{j=1}^{m}I^{j}_{t}<p_{N}/4)\leq\exp\left(-p_{N}^{2}m/8\right)$. Recall that this result is obtained assuming that $X_{1},\ldots,X_{T}$ are fixed. We take the union bound over all possible sequences of arms. However, a naïve union bound multiplies $K^{T}$ to the failure probability, which leads to an undesirable result of $m$ scaling linearly with $T$. We present the following proposition inspired by an observation from Lu and Van Roy [16] that a permutation of selected arms can be regarded as equivalent. We note that the strong result of Lemma 6 in Lu and Van Roy [16] is not applicable to our setting since we do not assume that $\{\eta_{t}\}_{t=1}^{T}$ is distributed identically nor independently. Although we present all the main ideas to support Proposition 1 in this section, there may be a few points that readers find require further justification. Due to limited space, we provide a full, rigorous justification of Proposition 1 in Appendix F, where we present a different perspective on the sampling of perturbation.
Proposition 1.
@@
The key observation in Proposition 1 is that the perturbation vector consists of i.i.d. components, hence its distribution is invariant under independent permutations. Therefore, the distributions of $\widetilde{\theta}_{t-1}^{j}$ and $U_{t-1}^{\top}{\mathbf{Z}}_{t}^{j}$ remain invariant under the permutation of selected arms. Although the sequence of arms and the perturbation vector are not independent as a whole, the permutation that sorts the selected arms preserves the distribution of $\mathbf{Z}_{t}$ since $X_{t}$ and $Z_{t}$ are independent for all $t\in[T]$. The number of equivalence classes up to permutation over sequences of arms with lengths at most $T-1$ is less than $T^{K}$, since each arm can be selected 00 to $T-1$ times inclusively. Therefore, we take the union bound over the $T^{K}$ sequences of arms and attain $\mathcal{E}_{2}^{*}$.
@@ 6 Ensemble Sampling and Perturbed-History Exploration
LinPHE is linear ensemble sampling with $T$ models: Conversely, since LinPHE can be regarded as linear ensemble sampling with $T$ models, it is possible to derive a regret bound of LinPHE by following the proof of Theorem 1. We present Corollary 1, which states a regret bound of $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for LinPHE. Note that in this case, the regret bound is independent of $K$.
@@
Kveton et al. [12] provide a $\widetilde{\mathcal{O}}(d\sqrt{T\log K})$ regret bound when the number of arms is finite. We prove that LinPHE achieves a $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret bound independent of the number of arms, using our analysis framework. It is widely observed that assuming the size of the arm set to be $K$ may lead to an interchanging of a $\sqrt{d}$ factor with a $\sqrt{\log K}$ factor in the regret bound [4], although attaining such reduction may not always be done in a trivial manner [5, 8, 6]. It it worth noting that $\widetilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret bound was previously achieved for LinPHE [janz2024exploration]. Our focus is not merely on proving another regret bound for LinPHE, but rather on highlighting the close relationship between linear ensemble sampling and LinPHE, which, to our knowledge, has been overlooked in the literature.
The proof of Corollary 1 follows the proof of Theorem 1. Note that the latter part of the proof of Theorem 1 focuses on decoupling the dependency between $\{X_{t}\}_{t=1}^{T}$ and $\mathbf{Z}_{t}$. However, in the case of LinPHE, they are already independent since the perturbation sequence is freshly sampled at every time step, enabling a more elegant and concise proof. Especially, as it skips the parts that require the number of arms to be finite, for instance the use of Proposition 1, Corollary 1 holds even when the number of arms is infinite. The whole proof is presented in Appendix G.
@@ Appendix A Notations
We summarize the notations in this paper in Table 2 and Table 3.
@@
$\mathcal{E}$ High-probability event
@@ Appendix B Proof of Theorem 2
As mentioned in the discussion of Theorem 2, we utilize Markov’s inequality and the random variable of interest is $X_{t}^{\top}\theta_{t}\mathds{1}\left\{\mathcal{E}_{1,t}\cap\mathcal{E}_{2,t}\right\}$. However, this random variable may not be non-negative, precluding the use of Markov’s inequality. The following lemma shows that adding an appropriate term guarantees its non-negativity.
@@
We lower-bound the probability on the left-hand side utilizing condition (3). Suppose the event of interest in condition (3), namely $((x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t})\lor\mathcal{E}_{2,t}^{\mathsf{C}}$, holds. Under the event, either $\mathcal{E}_{2,t}^{\mathsf{C}}$ or $(x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t}\land\mathcal{E}_{2,t}$ holds. Under $\mathcal{E}_{2,t}^{\mathsf{C}}$, $g_{t}(\theta^{*},\mathcal{E}_{2,t})\leq g_{t}\left(\theta_{t},\mathcal{E}_{1,t}\cap\mathcal{E}_{2,t}\right)$ becomes $0\leq 0$, which trivially holds. Otherwise, we have $(x^{*\top}\theta^{*}\leq X_{t}^{\top}\theta_{t})\land\mathcal{E}_{1,t}\land\mathcal{E}_{2,t}$. Under this event, we have that
@@
The second sum is bounded by the Azuma-Hoeffding inequality. Note that $0\leq\left\|X_{t}\right\|_{V_{t-1}^{-1}}\leq\sqrt{\lambda_{\max}(V_{t-1}^{-1})}\left\|X_{t}\right\|_{2}\leq\frac{1}{\sqrt{\lambda}}$, where $\lambda_{\max}(\cdot)$ denotes the maximum eigenvalue. By Lemma 12, with probability at least $1-\delta$, it holds that